1 mole of N2 and 4 mole of H2 are allowed to react in a vessel and after reaction H2O is added, the aqua solution required one mole of HCL of neutralization then mole fraction of H2 in the remaining gaseous mixture after reaction is-
A 1/6
B 5/6
C 1/3
D none of these

Grade:12

4 Answers

Arun
5 years ago
Dear Ayush
 
I am getting my answer as 5/6. i.e. option B.
 
Hope it helps
Regards
Arun (askIITians forum expert)
Ayush Mittal
5 years ago
Can You show the concept you applied and the complete steps involved in the solving the answer.
please respond as soon as possible.
Arun
5 years ago
Only 1 mole of NH3 gas is formed
 
Solution requires 1 mole of HCl and hence only 1 mole of NH3 is required.
 
HCl     +     NH3     ----->   NH4Cl
1 mole       1 mole
 
For NH3 formation
                        N2 +       3 H2 ------->      2 NH3
Initial moles    1 .          4.                         0
Final mole      1-x          4 -3x                    2x
 
2x = 1
 
Hence x = 0.5
 
Mole fraction of H2 gas in final mixtire
 
 = 2.5 /(0.5 + 2.5) = 2.5 /3 = 5/6
Ayush Mittal
5 years ago
Thank you iitian for my help. It will surely prove a great help to me. Thank you for your effort .
Regards .